OpenStudy (anonymous):

Find the volume of the cylinder to the nearest cubic foot. Use a calculator. A. 236 ft3 B. 942 ft3 C. 75 ft3 D. 251 ft3

OpenStudy (nurali):

volume of the cylinder = (pi)*r^2*h = (pi)(25)(3) = 236 ft^3

OpenStudy (kropot72):

The dimensions of the cylinder are needed to work out the volume. Can you please give the radius (or diameter) and the length (or height).

OpenStudy (anonymous):



OpenStudy (kropot72):

It looks like @Nurali already knew the question details.

OpenStudy (anonymous):

yeah she/he always right